# Building sets, Chow rings, and their Hilbert series

### v1

@@ 3. Hilbert series for arbitrary building sets
The main goal of this section is to extend [FMSV24, Theorem 1.4] to general building sets. To this end, we will work in the incidence algebra of the lattice of flats of a polymatroid. Specifically, if $\mathcal{L}(\mathsf{M})$ is the lattice of flats of a polymatroid, the incidence algebra $\mathcal{I}$ of $\mathcal{L}(\mathsf{M})$ is the free $\mathbb{Z}[x]$-module spanned by all closed intervals of $\mathcal{L}(\mathsf{M})$. The product of two elements $a,b\in\mathcal{I}$ is given by the convolution
@@
for every choice of flats $F\subseteq G$, where $a_{FG}$ denotes the polynomial that $a$ associates to the closed interval $[F,G]$. The two-sided identity in this algebra is the element $\delta\in\mathcal{I}$ defined by $\delta_{FG}:=1$ for $F=G$ and 00 otherwise. Another very useful object that we will use repeatedly is the so-called zeta function $\zeta\in\mathcal{I}$, which is defined by $\zeta_{FG}:=1$ for every $F\subseteq G$. We will assume that the reader is familiar with the basics of incidence algebras, and we refer to [Sta12, Chapter 3] for undefined terminology.
@@ 3.1. $\mathcal{G}$-reduced characteristic polynomials
If $\mathsf{M}$ is nonempty and $\mathcal{G}=\mathcal{G}_{\max}$, then $|f(\mathcal{G})|=|\{E\}|=1$, so the $\mathcal{G}_{\max}$-reduced characteristic polynomial agrees with the usual reduced characteristic polynomial $\overline{\chi}_{\mathsf{M}}(x):=\chi_{\mathsf{M}}(x)/(x-1)$. Since a building set $\mathcal{G}$ on $\mathcal{L}(\mathsf{M})$ induces building sets on each interval $[F,G]\subseteq\mathcal{L}(\mathsf{M})$, we can define the following key object in the incidence algebra of $\mathcal{L}(\mathsf{M})$.
@@
The unusual choice for how to define the $\mathcal{G}$-reduced characteristic polynomial of a singleton interval is important, and it is motivated by a similar choice made in [FMSV24] of defining the usual reduced characteristic polynomial of an empty matroid as $-1$. Before proving our main results, we need a further ingredient.
@@
It follows from the definitions that for every pair of flats $F\subseteq G$ of $\mathsf{M}$, we have $f(\mathcal{G}|_{G}/F)=f(\mathcal{G}|_{G})\setminus f(\mathcal{G}|_{F})$. Let us set $f(\mathcal{G}|_{G})=\{G_{1},G_{2},\dots,G_{k}\}$.
$\left(\zeta\cdot\overline{\chi}^{\mathcal{G}}\right)_{FG}$ $=-\sum_{F\subseteq H\subseteq G}\frac{\chi_{HG}(x)}{(1-x)^{\lvert f(\mathcal{G}|_{G}/H)\rvert}}$
$=-\sum_{F_{i}\in[F,G_{i}]}\prod_{\begin{subarray}{c}j=1\\ F_{j}\neq G_{j}\end{subarray}}^{k}\frac{\chi_{F_{j}G_{j}}(x)}{1-x}$
$=-\sum_{F_{i}\in[F,G_{i}]}\prod_{j=1}^{k}\left(-\overline{\chi}_{F_{j}G_{j}}(x)\right)$
$=(-1)^{|f(\mathcal{G}|_{G}/F)-1|}\sum_{F_{i}\in[F,G_{i}]}\prod_{j=1}^{k}\overline{\chi}_{F_{j}G_{j}}(x)$
$=(-1)^{|f(\mathcal{G}|_{G}/F)-1|}\prod_{G_{j}\supseteq F}\sum_{F_{j}\in[F,G_{j}]}\overline{\chi}_{FG_{j}}(x)$
$=(-1)^{|f(\mathcal{G}|_{G}/F)-1|}\prod_{G_{j}\supseteq F}(x+x^{2}+\cdots+x^{\operatorname{rk}(G_{j})-\operatorname{rk}(F)-1})=\alpha^{\mathcal{G}}_{FG}.$
The sum in the second line is over all choices of flats $(F_{1},\ldots,F_{k})\in[F,G_{1}]\times\cdots\times[F,G_{k}]$, and the factor inside is set to be $1$ when $F\not\subseteq G_{i}$ for some $i$; this explains the appearance of exactly $|f(\mathcal{G}|_{G}/F)|$ minus signs in the fourth line. Meanwhile, in the last line we used [FMSV24, Lemma 2.5]. ∎
@@ 3.2. Recursions for Hilbert series of Chow rings
[…]
@@
The range of $\varphi_{G}$ is the set of all FY-monomials associated to a nested set $N$ such that $f(\mathcal{G}|_{G})\subseteq\min(N)$. The sets $\operatorname{Im}\varphi_{G}$ cover $\operatorname{FY}(\mathsf{M},\mathcal{G})$ and an FY-monomial $m(N,k)=\prod_{F\in N}x_{F}^{k(F)}$ is contained in $2^{\lvert\min(N)\rvert}$ sets corresponding to the element $\vee T$ for any $T\subseteq\min(N)$. Therefore,
@@ 3.3. Explicit formulas for the Hilbert series
$\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{\varnothing\neq F\in\mathcal{L}(\mathsf{M})}\frac{-\chi_{\mathsf{M}|_{F}}(x)}{(1-x)^{|f(\mathcal{G}|_{F})|}}\cdot\mathrm{H}_{\mathsf{M}/F}^{\mathcal{G}/F}(x),$ (3)
@@
$\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{E\neq F\in\mathcal{L}(\mathsf{M})}\mathrm{H}_{\mathsf{M}|_{F}}^{\mathcal{G}|_{F}}(x)\cdot\frac{-\chi_{\mathsf{M}/F}(x)}{(1-x)^{\lvert f(\mathcal{G}/F)\rvert}}.$ (4)
@@
[…]
@@
Consider the loopless polymatroid $\mathsf{M}$ on $E=\{a,b,c\}$ described in [PP23, Section 7]. Its lattice of flats is depicted in Figure 2. Consider the building set $\mathcal{G}=\{a,b,c,abc\}$. One can enumerate the FY-monomials to see that
@@
Combining the identity in Theorem 3.6 with a general formula describing inverses in the incidence algebra as sums over chains (see [ER95, Lemma 5.3]) gives the result. ∎
@@
  • one flag of type $|3|4\subsetneq 12$ 3 | 4 ⊊ 1234,
@@ 3.4. A more efficient non-recursive formula
Following the terminology in [Cor25a, Definition 2.12], we say that a $\mathcal{G}$-nested set $S$ is spanning if the join of all the elements in $S$ is the ground set $E$ or, equivalently, if $S$ contains all the maximal elements of $\mathcal{G}$. We have the following formula for computing $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)$ in terms of spanning $\mathcal{G}$-nested sets.
@@
$\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{S}\prod_{F\in S}\overline{\chi}_{\mathsf{M}|_{F}/\sup(S,F)}(x),$ (5)
@@
$\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{E\neq F\in\mathcal{L}(\mathsf{M})}\left(\sum_{\begin{subarray}{c}S^{\prime}\text{ spanning }\\ \mathcal{G}|_{F}\text{-nested}\end{subarray}}\prod_{G\in S^{\prime}}\overline{\chi}_{\mathsf{M}|_{G}/\sup(S^{\prime},G)}\right)\overline{\chi}^{\mathcal{G}/F}_{\mathsf{M}/F}(x).$ (6)
@@ 3.5. The geometry of the formulas
Given a building set $\mathcal{G}$ containing $E$, let $\{G_{1},\dotsc,G_{m}\}$ be a total order on $\mathcal{G}$ that refines the order on $\mathcal{L}(\mathsf{M})$. The wonderful variety $\underline{W}_{L}^{\mathcal{G}}$ introduced in [DCP95] is the sequential blow-up of (the strict transforms of) the $\mathbb{P}L_{G_{i}}$, that is,
@@
Let $P_{X}(q)$ denote the virtual Poincaré polynomial of a variety $X$ over $\mathbb{C}$, characterized by the following two properties (see, for instance, [Ful93, Section 4.5]): (i) If $X$ is complete and smooth, then $P_{X}(q)=\sum_{i\geq 0}\dim H^{i}(X,\mathbb{Q})q^{i}$; and (ii) if $X=\bigsqcup_{j}Y_{j}$ is a decomposition of $X$ into finitely many locally closed subsets $Y_{j}$, then $P_{X}=\sum_{j}P_{Y_{j}}$. Note that $P_{\underline{W}_{L}^{\mathcal{G}}}(q)~{}=~{}\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(q^{2})$. Since $E\in\mathcal{G}$, the recursion (4) states
@@
  • The projection map $\pi\colon\underline{W}_{L}^{\mathcal{G}}\to\mathbb{P}L$ is a composition of blow-down maps, under which the strict transform of $\mathbb{P}L_{F}$ is $\underline{W}_{L^{F}}^{\mathcal{G}|_{F}}\times\underline{W}_{L_{F}}^{\mathcal{G}/F}$ [DCP95, Theorem 4.3], so that $\pi^{-1}(\mathbb{P}\mathring{L}_{F})\simeq\underline{W}_{L^{F}}^{\mathcal{G}|_{F}}\times\mathbb{P}\mathring{L}_{F}$.
@@
  • The virtual Poincaré polynomial $P_{\mathbb{P}\mathring{L}_{F}}(q)$ of $\mathbb{P}\mathring{L}_{F}$ is equal to the reduced characteristic polynomial $\overline{\chi}_{\mathsf{M}/F}(q^{2})$ (see for instance [Kat16, Section 7.2]).
4. Formulas for ¯$\overline{\mathcal{M}}_{0,n+1}$.
In this section we will apply our results to the case in which $\mathsf{M}=\mathsf{K}_{n}$ is the graphic matroid associated to the complete graph on $n$ vertices, and $\mathcal{G}=\mathcal{G}_{\min}$. These matroids are often called braid matroids since they admit realizations by the braid arrangements of type $A$. In this realization, the wonderful variety as recalled in Section 3.5 is the moduli space ¯$\overline{\mathcal{M}}_{0,n+1}$ of $(n+1)$-pointed stable rational curves [DCP95, Section 4.3]. The Poincaré polynomial $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$ of the cohomology ring of ¯$\overline{\mathcal{M}}_{0,n+1}$ therefore satisfies
@@
There have been several works on this Poincaré polynomial [Kee92, Man95, Get95, AMN24]. We apply our formulas from Section 3 to deduce a new formula for $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$ and recover previous results about $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$.
@@ 4.1. A formula in terms of partitions
Let us begin with an elementary but useful observation: the flats of $\mathsf{K}_{n}$ of rank $i$ are in bijection with the partitions of $[n]=\{1,\dotsc,n\}$ into $n-i$ nonempty subsets. The bijection is given by assigning to the partition $S_{1}\sqcup\dotsb\sqcup S_{n-i}=[n]$ the set of edges in the union $\mathsf{K}_{S_{1}}\sqcup\dotsb\sqcup\mathsf{K}_{S_{n-i}}$ of complete graphs. Thus, the connected flats of $\mathsf{K}_{n}$ correspond to partitions of $[n]$ in which there is at most one part that is not a singleton. The meet (resp. join) on the lattice of flats $\mathcal{L}(\mathsf{K}_{n})$ can be interpreted as the coarsest common refinement (resp. finest common coarsening) of partitions of $[n]$. For braid matroids, because a minor $\mathsf{M}|_{G}/F$ by flats $F$ and $G$ is again always a direct sum of smaller braid matroids (up to simplification of some parallel elements), one deduces that $\mathcal{G}_{\min}|_{G}/F$ is the minimal building set for the minor $\mathsf{M}|_{G}/F$. From now on, we write $\lambda\vdash[a]$ to denote a set partition $\lambda$ of the set $\{1,\ldots,a\}$.
To apply Theorem 1.3, we need a description of (spanning) $\mathcal{G}_{\min}$-nested sets of $\mathsf{K}_{n}$; this is accompished by the next proposition. This result has been rediscovered multiple times in the literature; see, e.g., Erdös–Székely [ES89, Theorem 1] or, for a more explicit explanation, [Gai15, Theorem 2.1] by Gaiffi. We include a formulation more adequate for our purposes, and a proof for the sake of completeness.
@@
There is a bijection between the set of spanning $\mathcal{G}_{\min}$-nested sets of cardinality $m$ in $\mathsf{K}_{n}$ and the collection of set partitions of $[n+m-1]$ into $m$ parts, none of which has size one. Moreover, for each spanning $\mathcal{G}_{\min}$-nested set, the sizes of the parts of the partition are determined as follows: each $F\in S$ corresponds to a part of size $k$ where $\mathsf{K}_{k}$ is the simplification of the minor $\mathsf{K}_{n}|_{F}/\operatorname{sup}(F,S)$.
@@
Let us construct an explicit bijection. Consider the set $A=[n]\cup\{w_{1},\ldots,w_{m-1}\}$ consisting of $n+m-1$ elements, i.e., the vertices of $\mathsf{K}_{n}$ and $m-1$ additional elements. Consider a spanning $\mathcal{G}$-nested set $S$, and linearly order the flats of $S$ as $G_{1},\ldots,G_{m}$ first by rank, then, among elements of the same rank, by their minimal element. Reading the list in this order, we build a partition of $A$ as follows:
@@
  • If $|S|>1$, we set a part in our partition to be $G_{1}$. To get the remaining parts, proceed as follows. First, replace all the flats $G\in S$ such that $G\supseteq G_{1}$ by $(G\setminus G_{1})\cup\{w_{1}\}$. Then, construct a partition of size $m-1$ of $([n]\setminus G_{1})\cup\{w_{2},\ldots,w_{m-1}\}$: inductively, this corresponds to a $\mathcal{G}_{\min}$-nested set of cardinality $m-1$ in $\mathsf{K}_{n-|G_{1}|}$.
@@
Now we state a technical lemma that will allow us to rewrite the formula in the previous corollary as a sum over set partitions of $[n-1]$ instead of $[n+m-1]$ for varying $m$.
@@
The Poincaré polynomial of ¯$\overline{\mathcal{M}}_{0,n+1}$ is

### v2

@@ 3. Hilbert series for arbitrary building sets
The main goal of this section is to extend [21, Theorem 1.4] to general building sets. To this end, we will work in the incidence algebra of the lattice of flats of a polymatroid. Specifically, if $\mathcal{L}(\mathsf{M})$ is the lattice of flats of a polymatroid, the incidence algebra $\mathcal{I}$ of $\mathcal{L}(\mathsf{M})$ is the free $\mathbb{Z}[x]$-module spanned by all closed intervals of $\mathcal{L}(\mathsf{M})$. The product of two elements $a,b\in\mathcal{I}$ is given by the convolution
@@
for every choice of flats $F\subseteq G$, where $a_{FG}$ denotes the polynomial that $a$ associates to the closed interval $[F,G]$. The two-sided identity in this algebra is the element $\delta\in\mathcal{I}$ defined by $\delta_{FG}:=1$ for $F=G$ and 0 otherwise. Another very useful object that we will use repeatedly is the so-called zeta function $\zeta\in\mathcal{I}$, which is defined by $\zeta_{FG}:=1$ for every $F\subseteq G$. We will assume that the reader is familiar with the basics of incidence algebras, and we refer to [42, Chapter 3] for undefined terminology.
@@ 3.1. $\mathcal{G}$-reduced characteristic polynomials
If $\mathsf{M}$ is nonempty and $\mathcal{G}=\mathcal{G}_{\max}$, then $|f(\mathcal{G})|=|\{E\}|=1$, so the $\mathcal{G}_{\max}$-reduced characteristic polynomial agrees with the usual reduced characteristic polynomial $\overline{\chi}_{\mathsf{M}}(x):=\chi_{\mathsf{M}}(x)/(x-1)$. Since a building set $\mathcal{G}$ on $\mathcal{L}(\mathsf{M})$ induces building sets on each interval $F,G]\subseteq\mathcal{L}(\mathsf{M})$, we can define the following key object in the incidence algebra of $\mathcal{L}(\mathsf{M})$.
@@
The unusual choice for how to define the $\mathcal{G}$-reduced characteristic polynomial of a singleton interval is important, and it is motivated by a similar choice made in [21] of defining the usual reduced characteristic polynomial of an empty matroid as 1-1. Before proving our main results, we need a further ingredient.
@@
For the sake of simplicity, and without loss of generality, we will prove that $\left(\zeta\cdot\overline{\chi}^{\mathcal{G}}\right)_{\varnothing E}=\alpha^{\mathcal{G}}_{\varnothing E}$. The general case follows from this one. Let us set $f(\mathcal{G})=\{G_{1},G_{2},\dots,G_{k}\}$. We have the following chain of equalities:
$\left(\zeta\cdot\overline{\chi}^{\mathcal{G}}\right)_{\varnothing E}$ $=-\sum_{H\in\mathcal{L}(\mathsf{M})}\frac{\chi_{HE}(x)}{(1-x)^{\lvert f(\mathcal{G}/H)\rvert}}$
$=-\sum_{F_{i}\in[\varnothing,G_{i}]}\prod_{\begin{subarray}{c}j=1\\ F_{j}\neq G_{j}\end{subarray}}^{k}\frac{\chi_{F_{j}G_{j}}(x)}{1-x}$
$=-\sum_{F_{i}\in[\varnothing,G_{i}]}\prod_{j=1}^{k}\left(-\overline{\chi}_{F_{j}G_{j}}(x)\right)$
$=(-1)^{k-1}\sum_{F_{i}\in[\varnothing,G_{i}]}\prod_{j=1}^{k}\overline{\chi}_{F_{j}G_{j}}(x)$
$=(-1)^{k-1}\prod_{j=1}^{k}\sum_{F_{j}\in[\varnothing,G_{j}]}\overline{\chi}_{F_{j}G_{j}}(x)$
$=(-1)^{k-1}\prod_{j=1}^{k}(x+x^{2}+\cdots+x^{\operatorname{rk}(G_{j})-1})=\alpha^{\mathcal{G}}_{\varnothing E}.$
The sum in the second line is over all choices of flats $(F_{1},\ldots,F_{k})\in[\varnothing,G_{1}]\times\cdots\times[\varnothing,G_{k}]$. The equality in the fifth line is obtained by exchanging the order of the sums and the products. Finally, in the last line we used [21, Lemma 2.5] and the fact that $k=|f(\mathcal{G})|$. ∎
@@ 3.2. Recursions for Hilbert series of Chow rings
and it counts (without considering the signs) the possible choices of an element of $A(G)$ and of an FY-monomial in $[G,E]$. For each $F\not\subseteq G$ define $\psi(F)\coloneqq\sup\left(f(\mathcal{G}|_{F})\setminus f(\mathcal{G}|_{G})\right)$ and extend $\psi$ to a map $N\mapsto\{\psi(F)\mid F\in N\}$. It is a straightforward check that this map is a bijection between $\mathcal{G}/G$-nested sets and $\mathcal{G}$-nested sets with no elements smaller or equal to $G$. Let $\operatorname{FY}(\mathsf{M},\mathcal{G})$ be the set of FY-monomials associated to $\mathsf{M}$ and $\mathcal{G}$, and define an injection (cf. [36, Lemma 4.14])
@@
The range of $\varphi_{G}$ is the set of all FY-monomials associated to a nested set $N$ such that $f(\mathcal{G}|_{G})\subseteq\min(N)$. The sets $\operatorname{Im}\varphi_{G}$ cover $\operatorname{FY}(\mathsf{M},\mathcal{G})$ and an FY-monomial $m(N,k)=\prod_{F\in N}x_{F}^{k(F)}$ is contained in $2^{\lvert\min(N)\rvert}$ sets corresponding to the element $\sup T$ for any $T\subseteq\min(N)$. Therefore,
@@ 3.3. Explicit formulas for the Hilbert series
(3) $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{\varnothing\neq F\in\mathcal{L}(\mathsf{M})}\frac{-\chi_{\mathsf{M}|_{F}}(x)}{(1-x)^{|f(\mathcal{G}|_{F})|}}\cdot\mathrm{H}_{\mathsf{M}/F}^{\mathcal{G}/F}(x),$
@@
(4) $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{E\neq F\in\mathcal{L}(\mathsf{M})}\mathrm{H}_{\mathsf{M}|_{F}}^{\mathcal{G}|_{F}}(x)\cdot\frac{-\chi_{\mathsf{M}/F}(x)}{(1-x)^{\lvert f(\mathcal{G}/F)\rvert}}$.
@@
$abc\varnothingaabbccabab$$\scriptstyle 2$$\scriptstyle 2$$\scriptstyle 2$$\scriptstyle 2$$\scriptstyle 1$$\scriptstyle 4$$\scriptstyle 1$
@@
Consider the loopless polymatroid $\mathsf{M}$ on $E=\{a,b,c\}$ described in [36, Section 7]. Its lattice of flats is depicted in Figure 2. Consider the building set $\mathcal{G}=\{a,b,c,abc\}$. One can enumerate the FY-monomials to see that
@@
Combining the identity in Theorem 3.6 with a general formula describing inverses in the incidence algebra as sums over chains (see [17, Lemma 5.3]) gives the result. ∎
@@
  • one flag of type $1|2|3|4\subsetneq 1234$,
@@ 3.4. A more efficient non-recursive formula
Following the terminology in [12, Definition 2.12], we say that a $\mathcal{G}$-nested set $S$ is spanning if the join of all the elements in $S$ is the ground set $E$ or, equivalently, if $S$ contains all the maximal elements of $\mathcal{G}$. We have the following formula for computing $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)$ in terms of spanning $\mathcal{G}$-nested sets.
@@
(5) $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{S}\prod_{F\in S}\overline{\chi}_{\mathsf{M}|_{F}/\sup(S,F)}(x),$
@@
(6) $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{E\neq F\in\mathcal{L}(\mathsf{M})}\left(\sum_{\begin{subarray}{c}S^{\prime}\text{ spanning }\\ \mathcal{G}|_{F}\text{-nested}\end{subarray}}\prod_{G\in S^{\prime}}\overline{\chi}_{\mathsf{M}|_{G}/\sup(S^{\prime},G)}\right)\overline{\chi}^{\mathcal{G}/F}_{\mathsf{M}/F}(x).$
@@ 3.5. The geometry of the formulas
Given a building set $\mathcal{G}$ containing $E$, let $\{G_{1},\dotsc,G_{m}\}$ be a total order on $\mathcal{G}$ that refines the order on $\mathcal{L}(\mathsf{M})$. The wonderful variety $\underline{W}_{L}^{\mathcal{G}}$ introduced in [15] is the sequential blow-up of (the strict transforms of) the $\mathbb{P}L_{G_{i}}$, that is,
@@
Let $P_{X}(q)$ denote the virtual Poincaré polynomial of a variety $X$ over $\mathbb{C}$, characterized by the following two properties (see, for instance, [24, Section 4.5]): (i) If $X$ is complete and smooth, then $P_{X}(q)=\sum_{i\geq 0}\dim H^{i}(X,\mathbb{Q})q^{i}$; and (ii) if $X=\bigsqcup_{j}Y_{j}$ is a decomposition of $X$ into finitely many locally closed subsets $Y_{j}$, then $P_{X}=\sum_{j}P_{Y_{j}}$. Note that $P_{\underline{W}_{L}^{\mathcal{G}}}(q)~=~\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(q^{2})$. Since $E\in\mathcal{G}$, the recursion (4) states
@@
  • The projection map $\pi\colon\underline{W}_{L}^{\mathcal{G}}\to\mathbb{P}L$ is a composition of blow-down maps, under which the strict transform of $\mathbb{P}L_{F}$ is $\underline{W}_{L^{F}}^{\mathcal{G}|_{F}}\times\underline{W}_{L_{F}}^{\mathcal{G}/F}$ [15, Theorem 4.3], so that $\pi^{-1}(\mathbb{P}\mathring{L}_{F})\simeq\underline{W}_{L^{F}}^{\mathcal{G}|_{F}}\times\mathbb{P}\mathring{L}_{F}$.
@@
  • The virtual Poincaré polynomial $P_{\mathbb{P}\mathring{L}_{F}}(q)$ of $\mathbb{P}\mathring{L}_{F}$ is equal to the reduced characteristic polynomial $\overline{\chi}_{\mathsf{M}/F}(q^{2})$ (see, for instance, [30, Section 7.2]).
4. Formulas for $\overline{\mathcal{M}}_{0,n+1}$.
In this section we will apply our results to the case in which $\mathsf{M}=\mathsf{K}_{n}$ is the graphic matroid associated to the complete graph on $n$ vertices, and $\mathcal{G}=\mathcal{G}_{\min}$. These matroids are often called braid matroids since they admit realizations by the braid arrangements of type $A$. In this realization, the wonderful variety as recalled in Section 3.5 is the moduli space $\overline{\mathcal{M}}_{0,n+1}$ of $(n+1)$-pointed stable rational curves [15, Section 4.3]. The Poincaré polynomial $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$ of the cohomology ring of $\overline{\mathcal{M}}_{0,n+1}$ therefore satisfies
@@
There have been several works on this Poincaré polynomial [31, 33, 26, 3]. We apply our formulas from Section 3 to deduce a new formula for $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$ and recover previous results about $P_{\overline{\mathcal{M}}_{0,n+1}}(x)$.
@@ 4.1. A formula in terms of partitions
Let us begin with an elementary but useful observation: the flats of $\mathsf{K}_{n}$ of rank $i$ are in bijection with the partitions of $n]=\{1,\dotsc,n\}$ into nin-i nonempty subsets. The bijection is given by assigning to the partition $S_{1}\sqcup\dotsb\sqcup S_{n-i}=[n]$ the set of edges in the union $\mathsf{K}_{S_{1}}\sqcup\dotsb\sqcup\mathsf{K}_{S_{n-i}}$ of complete graphs. Thus, the connected flats of $\mathsf{K}_{n}$ correspond to partitions of $[n]$ in which there is at most one part that is not a singleton. The meet (resp. join) on the lattice of flats $\mathcal{L}(\mathsf{K}_{n})$ can be interpreted as the coarsest common refinement (resp. finest common coarsening) of partitions of $[n]$. For braid matroids, because a minor $\mathsf{M}|_{G}/F$ by flats $F$ and $G$ is again always a direct sum of smaller braid matroids (up to simplification of some parallel elements), one deduces that $\mathcal{G}_{\min}|_{G}/F$ is the minimal building set for the minor $\mathsf{M}|_{G}/F$. From now on, we write $\lambda\vdash[a]$ to denote a set partition $\lambda$ of the set $\{1,\ldots,a\}$.
To apply Theorem 1.3, we need a description of (spanning) $\mathcal{G}_{\min}$-nested sets of $\mathsf{K}_{n}$; this is accomplished by the next proposition. This result has been rediscovered multiple times in the literature; see, e.g., Erdös–Székely [18, Theorem 1] or, for a more explicit explanation, [25, Theorem 2.1] by Gaiffi. We include a formulation more adequate for our purposes, and a proof for the sake of completeness.
@@
There is a bijection between the set of spanning $\mathcal{G}_{\min}$-nested sets of cardinality $m$ in $\mathsf{K}_{n}$ and the collection of set partitions of [n+m1][n+m-1] into $m$ parts, none of which has size one. Moreover, for each spanning $\mathcal{G}_{\min}$-nested set, the sizes of the parts of the partition are determined as follows: each $F\in S$ corresponds to a part of size $k$ where $\mathsf{K}_{k}$ is the simplification of the minor $\mathsf{K}_{n}|_{F}/\operatorname{sup}(F,S)$.
@@
Let us construct an explicit bijection. Consider the set $A=[n]\cup\{w_{1},\ldots,w_{m-1}\}$ consisting of n+m1n+m-1 elements, i.e., the vertices of $\mathsf{K}_{n}$ and m1m-1 additional elements. Consider a spanning $\mathcal{G}$-nested set $S$, and linearly order the flats of $S$ as $G_{1},\ldots,G_{m}$ first by rank, then, among elements of the same rank, by their minimal element. Recall that a nested set of $\mathsf{K}_{n}$ having $G_{1}$ as a minimal element is the same as a nested set in $\mathsf{K}_{n}/G_{1}$. Reading the list in the order described before, we build a partition of $A$ as follows:
@@
  • If $|S|>1$, we set a part in our partition to be $G_{1}$. To get the remaining parts, proceed as follows. First, replace all the flats $G\in S$ such that $G\supseteq G_{1}$ by $(G\setminus G_{1})\cup\{w_{1}\}$. Then, construct a partition with m1m-1 parts of $([n]\setminus(G_{1}\cup w_{1}))\cup\{w_{2},\ldots,w_{m-1}\}$: inductively, this corresponds to a $\mathcal{G}_{\min}$-nested set of cardinality m1m-1 in $\mathsf{K}_{n-|G_{1}|}$, as nested sets with $G_{1}$ as their minimal elements are in bijection with nested sets of $G_{1},E]$.
@@
Now we state a technical lemma that will allow us to rewrite the formula in the previous corollary as a sum over set partitions of [n1][n-1] instead of [n+m1][n+m-1] for varying $m$.
@@
The Poincaré polynomial of $\overline{\mathcal{M}}_{0,n+1}$ is